# Controlling superradiant phase transition in quantum Rabi model

### v1

@@ V Phase diagram
$S(r_{p})=\exp[\frac{r_{p}}{2}(a^{2}-a^{\dagger 2})]$ (34)
on $\tilde{H}_{p}$ (p=N, SX, SP, SXPa and SXPb). The detailed form of $r_{p}$ is given in the Appendix B. Then the resulting Hamiltonian reads
@@
$E_{G}^{\text{SXPa}}=$ $\-\frac{\tilde{\omega}_{c}}{2}(1+\frac{4}{\mu^{2}})-\tilde{\omega}_{0}(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}),$
$E_{G}^{\text{SXPb}}=$ $\frac{\tilde{\omega}_{c}}{2}(1-\frac{4}{\lambda^{2}})-\tilde{\omega}_{0}(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}).$
We plot the dependence of the first and second derivatives of $E_{G}$ on $\lambda$ and $\mu$ versus $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ in Fig. 6. Note that, $\lambda$ and $\mu$ are dimensionless rotating and CR coupling strengths, respectively. As shown in Fig. 6(a) and 6(b), it is found that $\partial E_{G}/\partial\lambda$ ($\partial E_{G}/\partial\mu$) is discontinuous at $g_{r}=0\ (g_{cr}=0)\ \text{for}\ \left|g_{cr}/g_{C}\right|>2$ ($\left|g_{r}/g_{C}\right|>2$). This results denote the QPT from SX to SP is of first order. As shown in Figs. 6(c) and 6(d), we observe that $\partial^{2}E_{G}/\partial\lambda^{2}$ ($\partial^{2}E_{G}/\partial\mu^{2}$) is discontinuous at the boundaries $g_{r}=\left|\tilde{g}_{C}\right|$ and $g_{r}=0$ ($g_{cr}=0$) for $\left|g_{cr}/g_{C}\right|>2$ ($\left|g_{r}/g_{C}\right|>2$). Thus, the transition from the normal phase to superradiant phases is of second order. We also use the dashed and solid lines to indicate the first and second order QPT in Fig. 5, respectively.
@@ VI Manipulation of Quantum phases
In this section, we discuss how to manipulate the ground state by varying the modulation parameters. We plot the excitation energy $\omega$ as a function of $\xi$ at $\nu/\omega_{0}=0.68$ in Fig. 7. The subscript “p” that labels the specific phase in $\omega_{p}$ is omitted.
Figure 7: (Color online) The excitation energy $\omega$ versus the dimensionless modulation amplitude $\xi$. Parameters are $\eta=10^{2},\nu/\omega_{0}=0.68$ and $g/\omega_{0}=0.06$.
@@
The solid curve indicates that the system can enter different phases by adjusting the dimensionless modulation amplitude $\xi$. The QPT occurs at $\xi=1.10172$ and 2.59789, where $g_{r}/g_{C}=-1$, $g_{cr}/g_{C}=-1$. The ground state is in the normal phase when $1.101720<\xi<$ 1.10172 and $32.59789<\xi<$ 3, and the SX phase appears when $2.597891.10172<\xi<$ 2.59789. The system experiences the QPTs from the normal phase to SX phase at $\xi=1.10172$ and from SX phase to the normal phase at $\xi=2.59789$.
@@
To see the system how to cross the phase boundary under the chosen parameter in Fig. 7, we plot the parameterized curves of [$g_{r}(\xi)/g_{C}$, $g_{cr}(\xi)/g_{C}$] versus the modulation amplitude $\xi$ in Fig. 8(a). The parameterized curve evolves from the normal phase to SX phase then back to the normal phase as $\xi$ increases from 0 to 3. The system can be described by a standard QRM. In Fig. 8(b), we plot $\langle x\rangle$ versus the modulation amplitude $\xi$ for the same parameters as Fig. 8(a). Note that, $\langle p\rangle=0$ under the parameters in Fig. 8(a). As shown in Fig. 8(b), although the coupling strength is only in the strong-coupling regime with $g/\omega_{0}=0.06$, yet by the modulation, the QRM can reach the superradiant phase when $2.597891.10172<\xi<$ 2.59789. The normal phase appears when $1.101720<\xi<$ 1.10172 and $32.59789<\xi<$ 3. The results shown in Fig. 8 are consistent with those in Fig. 7. Thus, we can manipulate quantum phases on the QRM via the appropriate modulation with strong coupling.
@@
We mention that, the anisotropic QRM can be realized with our scheme, as shown in Fig. 2(b). However, to reach the classical field limit $\tilde{\eta}\rightarrow\infty$, only an effective standard QRM can be realized. This is because in the limit $\tilde{\eta}\rightarrow\infty$ and the condition (11), we obtain $m_{0}=n_{0}$, thus $g_{r}=g_{cr}$. This is an effective standard QRM. Despite this, we still provide a possible scheme to observe the superradiant phase of few-body quantum systems in the strong-coupling regime.
@@
In this paper, we consider negligible cavity dissipation with $\kappa/\omega_{0}<10^{-4},\ \kappa/\omega_{c}^{\prime}<10^{-2}$. When the cavity dissipation is non-negligible, it will produce a shift to the critical coupling strength, which becomes $g_{C}^{\text{diss}}=\frac{1}{2}\sqrt{\frac{\tilde{\omega}_{0}}{\tilde{\omega}_{c}}(\tilde{\omega}_{c}^{2}+\kappa^{2})}$. For a dissipation rate $\kappa=0.2\omega_{c}^{\prime}=0.002\omega_{0}$ and $\nu/\omega_{0}=0.68$, $\tilde{\omega}_{c}=0.01\omega_{0}$ and $g_{C}^{\text{diss}}\approx 1.02g_{C}$. The effective coupling strengths $g_{r}$ and $g_{cr}$ can be larger than $g_{C}^{\text{diss}}$, as shown in Fig. 2(c). Hence, it is still effective for our approach to manipulate the steady-state superradiant phase transition of the dissipative QRM.
VII two-photon $A^{2}$ terms
@@
In this section, we discuss the validity of omitting the two-photon $A^{2}$ terms. The two-photon $A^{2}$ terms given by Hamiltonian (5) can be written as
@@
The first term in $H_{A^{2}}$ has been absorbed into the Hamiltonian of the standard QRM by replacing the bosonic mode frequency $\omega_{c}$ with the modified frequency $\omega_{c}^{\prime}=\omega_{c}+2g_{A^{2}}$. The second term can be safely neglected with $g\ll\omega_{0},\ g_{A^{2}}\ll 2\omega_{c}^{\prime}$. To demonstrate this, in Fig. 9, we plot the ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g$ and the dimensionless coefficient $\chi$. For the dispersion case $\eta=10^{2}$, as shown in Fig. 9(a), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, as indicated by the dashed curve, $g$ and $\chi$ need to be ($g/\omega_{0}\sim 0.06,\ \chi\sim 0.0579$) or ($g/\omega_{0}\sim 0.0102,\ \chi\sim 2$). For the parameters we used in our discussions with $\eta=10^{2}$, $g/\omega_{0}=0.06$, and $\chi\leq 0.05$, this ratio is smaller than $1\%$, namely, $g_{A^{2}}/2\omega_{c}^{\prime}<0.01.$ Hence, we can neglect these terms in the dispersion limit. For the resonance case $\eta=1$, as shown in Fig. 9(b), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, we should have ($g/\omega_{0}\sim 0.06,\ \chi\sim 5.787$). Therefore, it is also reasonable to omit the $A^{2}$ terms for $g/\omega_{0}=0.06$ and $\chi<5$ for resonance case $\eta=1$. Thus we can omit the $A^{2}$ terms from the following discussions.
Figure 9: (Color online) The ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g/\omega_{0}$ and the dimensionless factor $\chi$. Parameters are (a) $\eta=10^{2}$ and (b) $\eta=1$.
VIII conclusion and discussions
@@
By applying a qubit frequency modulation to a standard QRM, we generate an anisotropic QRM whose effective couplings for rotating and counter-rotating terms can be adjusted over a wide range. We calculate the ground-state and excitation energy of this model in terms of the modulation parameters. The phase diagram of the anisotropic QRM in the whole coupling parameter space and the type of the QPT are also obtained. We show that the superradiant phase of the QRM can be observed in the strong-coupling regime with the frequency modulation. Furthermore, the two-photon $A^{2}$ terms, which could prevent the observation of the QPT in QRM Hwang2015Rabi , can be appropriately neglected in our scheme. We provide an experimentally accessible scheme to control the QPT in QRM. Our scheme is within current experimental technology trapped ions ; nuclear ; ZhengShiBiao ; 24.2 . This modulation scheme may also be applicable in other few-body quantum systems. Our work may pave the way for experimentally observing superradiant phase in other few-body quantum systems.
@@ Acknowledgements.
J.-F.H. is supported in part by the National Natural Science Foundation of China (Grant No. 12075083).
Appendix A Operators for superradiant phase
@@
The new Pauli operators we introduced are
@@
for SXPb phase, respectively. The form of unitary operator $U_{k}$ (k=SX, SP, SXPa, SXPb) for k phase reads
@@
$U_{\text{SXPa}}$ $=\exp[\frac{-ig_{c}}{{\sqrt{2}\Omega_{\text{SXPa}}}}(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\tau_{y}^{{}_{\text{SXPa}}}x-\tau_{x}^{{}_{\text{SXPa}}}p)],$
$U_{\text{SXPb}}$ $=\exp[\frac{-ig_{r}}{{\sqrt{2}\Omega_{\text{SXPb}}}}(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\tau_{y}^{{}_{\text{SXPb}}}x+\tau_{x}^{{}_{\text{SXPb}}}p)].$
Appendix B Squeezing parameters

### v2

@@ V Phase diagram
$S(r_{p})=\exp\left[\frac{r_{p}}{2}(a^{2}-a^{\dagger 2})\right]$ (35)
on $\tilde{H}_{p}$ ($p$=N, SX, SP, SXPa, and SXPb). The detailed form of $r_{p}$ is given in the Appendix C. Then the resulting Hamiltonian reads
@@
$E_{G}^{\text{SXPa}}=$ $-\frac{\tilde{\omega}_{c}}{2}\left(1+\frac{4}{\mu^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}\right),$
$E_{G}^{\text{SXPb}}=$ $-\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{4}{\lambda^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}\right).$
We plot the dependence of the first and second derivatives of $E_{G}$ on $\lambda$ and $\mu$ versus $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ in Fig. 6. Note that, $\lambda$ and $\mu$ are dimensionless rotating and CR coupling strengths, respectively. As shown in Fig. 6(a) and 6(b), it is found that $\partial E_{G}/\partial\lambda$ ($\partial E_{G}/\partial\mu$) is discontinuous at $g_{r}=0\ (g_{cr}=0)\ \text{for}\ \left|g_{cr}/g_{C}\right|>2$ ($\left|g_{r}/g_{C}\right|>2$). This result denotes the QPT from the SX phase to the SP phase is of first order. As shown in Figs. 6(c) and 6(d), we observe that $\partial^{2}E_{G}/\partial\lambda^{2}$ ($\partial^{2}E_{G}/\partial\mu^{2}$) is discontinuous at the boundaries $g_{r}=\left|\tilde{g}_{C}\right|$ and $g_{r}=0$ ($g_{cr}=0$) for $\left|g_{cr}/g_{C}\right|>2$ ($\left|g_{r}/g_{C}\right|>2$). Thus, the transition from the normal phase to superradiant phases is of second order. We also use the dashed and solid lines to indicate the first- and second-order QPT in Fig. 5, respectively.
@@ VI Manipulation of Quantum phases
In this section, we discuss how to manipulate the ground state by varying the modulation parameters. We plot the excitation energy $\omega$ as a function of $\xi$ at $\nu/\omega_{0}=0.68$ in Fig. 7. The subscript $p$ that labels the specific phase in $\omega_{\text{p}}$ is omitted. The solid curve indicates that the system can enter different phases by adjusting the dimensionless modulation amplitude $\xi$. The QPT occurs at $\xi=1.102$ and $2.598$, where $g_{r}/g_{C}=-1$, $g_{cr}/g_{C}=-1$. The ground state is in the normal phase when $1.1020<\xi<$ 1.102 and $32.598<\xi<$ 3, and the SX phase appears when $2.5981.102<\xi<$ 2.598. The system experiences the QPTs from the normal phase to the SX phase at $\xi=1.102$ and from the SX phase to the normal phase at $\xi=2.598$.
Figure 9: (Color online) The ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g/\omega_{0}$ and the dimensionless factor $\chi$. Parameters are (a) $\omega_{0}/\omega_{c}=10^{2}$ and (b) $\omega_{0}/\omega_{c}=1$.
@@
To see the system how crosses the phase boundary under the chosen parameter in Fig. 7, we plot the parameterized curves of [$g_{r}(\xi)/g_{C}$, $g_{cr}(\xi)/g_{C}$] versus the modulation amplitude $\xi$ in Fig. 8(a). The parameterized curve evolves from the normal phase to the SX phase and then back to the normal phase as $\xi$ increases from 0 to 3. The system can be described by a standard QRM. In Fig. 8(b), we plot $\langle x\rangle$ versus the modulation amplitude $\xi$ for the same parameters as Fig. 8(a). Note that, $\langle p\rangle=0$ under the parameters in Fig. 8(a). As shown in Fig. 8(b), although the coupling strength is only in the strong-coupling regime with $g/\omega_{0}=0.06$, yet by the modulation, the QRM can reach the superradiant phase when $2.5981.102<\xi<$ 2.598. The normal phase appears when $1.1020<\xi<$ 1.102 and $32.598<\xi<$ 3. The results shown in Fig. 8 are consistent with those in Fig. 7. Thus, we can manipulate quantum phases on the QRM via the appropriate modulation with strong coupling.
@@
We mention that the anisotropic QRM can be realized with our scheme, as shown in Fig. 2(b). However, to reach the classical field limit $\tilde{\eta}\rightarrow\infty$, only an effective standard QRM can be realized. This is because in the limit $\tilde{\eta}\rightarrow\infty$ and the condition (11), we obtain $m_{0}=n_{0}$, thus $g_{r}=g_{cr}$. This is an effective standard QRM. Despite this, we still provide a possible scheme to observe the superradiant phase of few-body quantum systems in the strong-coupling regime.
@@
In this paper, we consider a negligible cavity dissipation with $\kappa/\omega_{0}<10^{-4}$ and $\kappa/\omega_{c}^{\prime}<10^{-2}$ Johansson ; RMP ; Probing . When the cavity dissipation is non-negligible, it will produce a shift to the critical coupling strength, which becomes $g_{C}^{\text{diss}}=\frac{1}{2}\sqrt{\frac{\tilde{\omega}_{0}}{\tilde{\omega}_{c}}(\tilde{\omega}_{c}^{2}+\kappa^{2})}$ dispassive QRM . For a dissipation rate $\kappa=0.2\omega_{c}^{\prime}=0.002\omega_{0}$ and $\nu/\omega_{0}=0.68$, $\tilde{\omega}_{c}=0.01\omega_{0}$ and $g_{C}^{\text{diss}}\approx 1.02g_{C}$. The effective coupling strengths $g_{r}$ and $g_{cr}$ can be larger than $g_{C}^{\text{diss}}$, as shown in Fig. 2(c). Hence, it is still effective for our approach to manipulate the steady-state superradiant phase transition of the dissipative QRM.
VII Analysis concerning the two-photon $A^{2}$ term
It is known that the existence of superradiant phase transition in the cavity-QED system is subject to a no-go theorem, where the squared term of the electromagnetic vector potential (i.e., the $A^{2}$ term) leads to the disappearance of the QPT in the Dicke model A2 term . This no-go theorem also applies to the QPT of the QRM Hwang2015Rabi . In the cavity-QED system, the coefficient $\chi$ in the $A^{2}$ term satisfies $\chi\geq 1$ due to Thomas-Reiche-Kuhn sum rule no-go theorm . Thus, the occurrence of the superradiant phase transition is still challenging in the cavity-QED system. However, in the circuit-QED system, $\chi$ could be smaller than 1. Then the superradiant phase transition is recovered with a shifted critical point caused by the $A^{2}$ term A2 term ; Hwang2015Rabi ; no-go theorm ; 1A2 ; 2A2 ; 3A2 .
@@ VII Analysis concerning the two-photon $A^{2}$ term
In this section, we discuss the validity of omitting the two-photon $A^{2}$ term. The two-photon $A^{2}$ term given by Hamiltonian (5) can be written as
@@
The first term in $H_{A^{2}}$ is absorbed into the Hamiltonian of the standard QRM by replacing the bosonic mode frequency $\omega_{c}$ with the modified frequency $\omega_{c}^{\prime}=\omega_{c}+2g_{A^{2}}$. The second term can be safely neglected with $g\ll\omega_{0},\ g_{A^{2}}\ll 2\omega_{c}^{\prime}$. To demonstrate this, in Fig. 9, we plot the ratio $g_{A^{2}}/2\omega_{c}^{\prime}$ versus the coupling strength $g$ and the dimensionless coefficient $\chi$. For the dispersion case $\omega_{0}/\omega_{c}=10^{2}$, as shown in Fig. 9(a), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, as indicated by the dashed curve, $g$ and $\chi$ need to be ($g/\omega_{0}\sim 0.06,\ \chi\sim 0.058$) or ($g/\omega_{0}\sim 0.01,\ \chi\sim 2$). For the parameters we use in our discussions with $\omega_{0}/\omega_{c}=10^{2}$, $g/\omega_{0}=0.06$, and $\chi\leq 0.05$, this ratio is smaller than $1\%$, namely, $g_{A^{2}}/2\omega_{c}^{\prime}<0.01.$ Hence, we can neglect the $A^{2}$ term in the dispersion limit under the chosen parameters. For the resonance case $\omega_{0}/\omega_{c}=1$, as shown in Fig. 9(b), to satisfy $g_{A^{2}}/2\omega_{c}^{\prime}=0.01$, we should have ($g/\omega_{0}\sim 0.06,\ \chi\sim 5.787$). Therefore, it is also reasonable to omit the $A^{2}$ term for $g/\omega_{0}=0.06$ and $\chi<5$ for the resonance case $\omega_{0}/\omega_{c}=1$. Thus we can omit the $A^{2}$ term by choosing appropriate parameters.
VIII Discussions on the experimental implementation
In this section, we present some discussions concerning the experimental implementation of our scheme. The key element for the experimental implementation of this scheme is to realize the QRM with the strong coupling between the qubit and bosonic field, as well as the qubit frequency modulation. Here, we present some detailed discussions concerning the realization of the two points. On the one hand, the QPT of the QRM in our scheme occurs when the coupling strength between the qubit and the field is only in the strong-coupling regime, for example $g/\omega_{0}=0.06$. This coupling strength can be realized in various experimental platforms 3A2 ; ZhengShiBiao ; trapped ions ; LiJian ; Probing ; D. Lv ; cqed . To present some analyses on the current experimental conditions in these systems, the related experimental parameters reported in the circuit-QED system ZhengShiBiao and the trapped-ion system trapped ions are listed in Table 2. The frequency ratio is $\eta\simeq 10$ and 25 for the circuit-QED system ZhengShiBiao and the trapped-ion system trapped ions , respectively, which provide the condition for the occurrence of the QPT in QRM.
On the other hand, the qubit frequency modulation is also accessible in several experimental platforms. For a superconducting circuit system, the transition frequency of an Xmon qubit, which couples to a resonator, can be periodically modulated by an ac magnetic flux  ZhengShiBiao . In a trapped-ion system, we can use a linear Pauli trap to confine an ion to simulate the QRM. A pair of counter-propagating Raman laser beams can be used to modulate the transition frequency of the qubit formed by hyperfine states trapped ions . As reported in LiJian , the frequency of the qubit is $\omega_{0}/2\pi=2.62$ GHz and the modulation amplitude is $\xi\nu/2=2\pi\times 250$ MHz. One can vary the modulation frequency $\nu/2\pi$ from 0 to 500 MHz, leading to $0\leq\nu/\omega_{0}\leq 0.19$ and $\xi>1$. Thus, the modulation parameters reported in experiments are of the same order of magnitude as those used in our scheme. We list the parameters for simulations of our scheme in Table 3. All the parameters are achievable in experiments trapped ions ; LiJian ; ZhengShiBiao ; D. Lv . In addition, we consider the negligible dissipation of the field and qubit. This assumption can also be supported by current experimental technology, based on the fact that the dimensionless decay rates of the field and qubit could be smaller than $10^{-3}$ in the above experimental platforms Johansson ; RMP ; Probing .
IX conclusion
@@ VIII Discussions on the experimental implementation
We proposed a scheme for generating an anisotropic QRM by applying a qubit frequency modulation to a standard QRM. Here the effective couplings for both the rotating and counter-rotating terms in the obtained anisotropic QRM can be adjusted over a wide range. We calculated the ground-state and excitation energy of this model in terms of the modulation parameters. The phase diagram of the anisotropic QRM in the entire coupling parameter space and the type of the QPT were also obtained. It was shown that the superradiant phase of the QRM can be observed in the strong-coupling regime with the frequency modulation. Furthermore, the two-photon $A^{2}$ term, which could prevent the observation of the QPT in QRM Hwang2015Rabi , can be appropriately neglected in our scheme. We also provided an experimentally accessible scheme to control the QPT in QRM. Our scheme is within current experimental technology trapped ions ; nuclear ; ZhengShiBiao ; 24.2 . This modulation scheme may also be applicable in other few-body quantum systems. Our work will pave the way for experimentally observing the superradiant phase in other few-body quantum systems.
@@ Acknowledgements.
J.-F.H. is supported in part by the National Natural Science Foundation of China (Grants No. 12075083 and No.12475016) and the Key Program of Xiangjiang-Laboratory in Hunan province, China (Grant No. XJ2302001). We would like to thank Jie-Qiao Liao for the helpful discussions during the reply to the referees’ reports.
Appendix A Evaluation of the validity of the rotating-wave approximation
To evaluate the validity of the rotating wave approximation used in Hamiltonian (9), we examine the time evolution of the fidelity $F(t)=\left|\left\langle\phi\left(t\right)\right|\left.\psi\left(t\right)\right\rangle\right|^{2}$ between the states $\left|\phi\left(t\right)\right\rangle$ and $\left|\psi\left(t\right)\right\rangle$, which are governed by the Hamiltonian (9) and Hamiltonian (12), respectively. Here, we choose the initial state $\left|\phi\left(0\right)\right\rangle=\left|\psi\left(0\right)\right\rangle=(1/\sqrt{2})\left(\left|g\right\rangle+\left|e\right\rangle\right)\left|\alpha\right\rangle$, where $\left|\alpha\right\rangle$ is a coherent state of the cavity field.
Table 2: Experimental parameters reported in the circuit-QED system ZhengShiBiao and trapped-ion system trapped ions : the transition frequency $\omega_{0}$ of the two-level system, the frequency $\omega_{c}$ of the cavity field, the coupling strength $g$, and the ratio $g/\omega_{0}$.
Reference Description $\omega_{0}/2\pi$ $\omega_{c}/2\pi$ $g/2\pi$ $g/\omega_{0}$
ZhengShiBiao Circuit QED $3.42-10.25$ MHz $0.34-1.03$ MHz 0.81 MHz $0.079-0.237$
trapped ions Trapped ion 50 kHz 2 kHz $0-7.1$ kHz $0-0.142$
Table 3: Analysis of the used parameters in our scheme corresponding to the experimental parameters given in Table 2.
Used parameters Circuit QED Trapped ion
[…]
Dimensionless modulation amplitude $\xi$ $\xi=0-3$- --
[…]
We plot the fidelity $F(t)$ versus $t$ for $\nu/\omega_{0}=0.49,\ 0.68,\text{and}\ 1$. As shown in Fig. 10, the fidelity can be higher than 0.9 when $\omega_{0}t/2\pi\leq 9$ for the used parameters. Specifically, for $\nu/\omega_{0}=0.68,$ which is used in our simulations, the fidelity is higher than 0.97 for $\omega_{0}t/2\pi\leq 10$. In particular, the fidelity can be further improved by choosing a suitably larger $\nu$. This result implies that the RWA is valid in our scheme under the chosen parameters.
Appendix B Operators for superradiant phase
@@
The new Pauli operators we introduce are
@@
for the SXPb phase, respectively.
The form of unitary operator $U_{\text{k}}$ ($k$=SX, SP, SXPa, SXPb) for $k$ phase reads
@@
$U_{\text{SXPa}}$ $=\exp\left[\frac{-ig_{c}}{{\sqrt{2}\Omega_{\text{SXPa}}}}\left(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPa}}}\tau_{y}^{{}_{\text{SXPa}}}x-\tau_{x}^{{}_{\text{SXPa}}}p\right)\right],$
$U_{\text{SXPb}}$ $=\exp\left[\frac{-ig_{r}}{{\sqrt{2}\Omega_{\text{SXPb}}}}\left(\frac{\tilde{\omega}_{0}}{\Omega_{\text{SXPb}}}\tau_{y}^{{}_{\text{SXPb}}}x+\tau_{x}^{{}_{\text{SXPb}}}p\right)\right].$
Figure 10: (Color online) Dynamics of the fidelity $F(t)$ for $\nu/\omega_{0}=0.49,\ 0.68,\text{and}\ 1$, respectively. The initial state of the system is $\left|\phi\left(0\right)\right\rangle=\left|\psi\left(0\right)\right\rangle=(1/\sqrt{2})\left(\left|g\right\rangle+\left|e\right\rangle\right)\left|\alpha\right\rangle$ with $\alpha=0.1$. We take the modulation amplitude $\xi=1.5$. Other parameters are the same as Fig. 7.
Appendix C Squeezing parameters